# Leptogenesis from magnetic helicity of gauged $\text{U}(1)_{B-L}$

### v1

@@ Abstract
If the $B-L$ symmetry is gauged with the addition of right-handed neutrinos, the standard model $B-L$ current is anomalous with respect to the $B-L$ gauge field itself. Then, the anomaly relation implies that the magnetic helicity of the $B-L$ gauge field is related to the standard model $B-L$ charges, although the whole universe is $B-L$ neutral with right-handed neutrinos. Based on this, we propose a new leptogenesis scenario with the gauged $B-L$ symmetry as follows. First, the magnetic helicity of the $B-L$ gauge field is generated, e.g., by the axion inflation, together with the standard model and right-handed neutrino $B-L$ charges, with the net $B-L$ charge kept zero. The $B-L$ charges in the standard model and right-handed neutrino sectors are then subject to washout effects from the interactions between them. After the washout effects decouple, the $B-L$ gauge symmetry is Higgsed and the magnetic helicity of the $B-L$ gauge field decays and generates $B-L$ charges in the both sector; thanks to the washout effects, we obtain a non-zero $B-L$ asymmetry. We show that the baryon asymmetry of the universe can be generated in this scenario, discussing the decay of the magnetic helicity of the $B-L$ gauge field and the interactions between the right-handed neutrinos and the standard model particles.
@@ I Introduction
$B-L$ charges are not washed out by the standard model process as the $B-L$ symmetry in the standard model is not anomalous with respect to $\text{SU}(3)_{C}\times\text{SU}(2)_{L}\times\text{U}(1)_{Y}$. This, however, seems to conclude that $B-L$ cannot be generated with the magnetic helicity, at first glance. Ref. Domcke et al. (2021, 2023) has circumvent the apparent contradiction by introducing the right-handed neutrinos, which explains the neutrino mass Workman et al. (2022) through the seesaw mechanism Minkowski (1977); Yanagida (1979); Gell-Mann et al. (1979). The focus of the work is on the conservation of fermion chiralities in the standard model; some of the standard model Yukawa couplings, such as those for $e,u$, and $d$, are small enough at high temperatures and their chiralities are effectively conserved in the early universe. The fermion chiralities are anomalous with respect to the standard model gauge groups and can be generated associated with the magnetic helicity. $B-L$ violating interactions of the right-handed neutrinos, which become effective only when they are sufficiently massive before decay, can generate a non-zero $B-L$ in the thermal bath with non-zero fermion chirality, even when the Sakharov conditions Sakharov (1967) are not simultaneously satisfied. Note that, however, this scenario depends on the flavor and chirality structure in the UV physics; many extensions of the standard model predict different flavor and chirality structures Isidori et al. (2010).
@@
In this work, we present a new avenue for the generation of the $B-L$ asymmetry from magnetic helicity. Previous works assume that the $B-L$ symmetry is not broken by the anomaly of the standard model gauge groups. However, this is not precise; the $B-L$ symmetry has the gravitational anomaly and the ’t Hooft anomaly Weinberg (2013). The former is considered in the gravi-leptogenesis scenario Alexander et al. (2006), but, as Ref. Alexander et al. (2006) itself admits, the amount of the gravitational wave would be too much if we assumed all the baryon asymmetry of the universe is generated with helical gravitational waves. See also Refs. Fischler and Paban (2007); Kamada et al. (2020).
@@
On the other hand, the latter possibility has not been explored much, to our knowledge***See Ref. Chao et al. (2024) for a recent study to generate baryon asymmetry through the $B-L$ ’t Hooft anomaly in axion inflation.. That is, once the $B-L$ is gauged with the addition of right-handed neutrinos, the standard model $B-L$ symmetry is anomalous with respect to the gauge field itself. Based on this, we propose a new mechanism for the generation of the baryon asymmetry of the universe as follows. The magnetic helicity of the $B-L$ gauge field is generated in the early universe, e.g., during the axion inflation, and the $B-L$ charges in the standard model sector are generated with the magnetic helicity. The $B-L$ charges in the right-handed neutrino sector are also generated with the magnetic helicity, and the total amount of the $B-L$ charges in the universe is zero. By the interactions between the right-handed neutrinos and the standard model particles, the $B-L$ charges in both sectors once tend to be equilibrated and relaxed to zero while keeping the magnetic helicity of the $B-L$ gauge field conserved. Then, as the $B-L$ gauge symmetry is broken at the scale $v_{B-L}$, the magnetic helicity of the $B-L$ gauge field decays and the difference between the $B-L$ charges in the standard model sector and right-handed neutrino sector arises once more. If the washout interaction for the $B-L$ charge in the standard model sector is decoupled before the decay of the magnetic helicity, the difference remains as the baryon number of the universe. As our scenario directly generates $B-L$ charges, we may generate enough baryon asymmetry even if the flavor symmetry is badly broken in the UV physics.
@@
The organization of the paper is following. In Sec. II, we review the co-generation of the magnetic helicity and chirality and discuss the decay of the former. In Sec. III, we assume the interaction between the right-handed neutrinos and the standard model particles is such that the standard model neutrinos are Majorana particles and discuss the evolution of the $B-L$ charges in the standard model sector. We find that the result depends on the UV structure of the theory and the phase transition of the $B-L$ gauge symmetry breaking. In Sec. IV, we assume the standard model neutrinos are Dirac particles and discuss the evolution of the $B-L$ charges in the standard model sector. We find that the baryon asymmetry of the universe can be easily generated in this case. Finally, Sec. V, is devoted to the discussion and conclusion.
@@ II Generation of the helicity and its decay
where $\vec{A}$ is the spatial component of the gauge field and $\vec{B}$ is the magnetic field. The magnetic helicity is a gauge invariant quantity if the magnetic field dumps off at the spatial infinity. In the following, we assume the $B-L$ symmetry is gauged and consider the magnetic helicity of the $B-L$ gauge field, $X_{\mu}$, namely, $\mathcal{H}_{X}$.
The magnetic helicity has a close relation to the chiral anomaly of the gauge field; the change of the magnetic helicity corresponds to the change of the chiral charge of the fermions. If we ignore the neutrino mass, the standard model $B-L$ current, $J^{\mu}_{B-L,\text{SM}}$, is anomalous with respect to the $B-L$ gauge fieldHere, we take the covariant form of the anomaly Bardeen and Zumino (1984).;
@@
where $g_{B-L}$ is the coupling constant of the $B-L$ gauge field, $X_{\mu\nu}$ is the field strength tensor of $X_{\mu}$, and $\tilde{X}^{\mu\nu}$ is the dual tensor of $X_{\mu\nu}$. Integrating the above equation over the spatial volume, we obtain
@@
where $N_{B-L,\text{SM}}$ is the $B-L$ charge of the standard model particles in the volume. Here, we assume that all fields drop off quickly enough at the spatial infinity. This equation shows that the change of the magnetic helicity of the $B-L$ gauge field corresponds to the change of the $B-L$ charge of the standard model particles.
In the following, we focus mainly on the dynamics of the $B-L$ gauge field in the context of axion inflation. In the subsequent sections, we will analyze the evolution of the $B-L$ asymmetry in the matter sector, taking into account the generation of $B-L$ helicity in axion inflation. For this purpose, we need to specify the model of the right-handed neutrinos.
@@ II.1 Generation of the helicity
The Lagrangian of the inflaton $\phi$ and the $B-L$ gauge field is given by
@@
where polarization vector ϵ(±)superscriptitalic-$\vec{\epsilon}^{\,(\pm)}$ satisfying $\vec{\epsilon}^{\,(\pm)}(\vec{k})\cdot\vec{k}=0,\ i\vec{k}\times\vec{\epsilon}^{\,(\pm)}(\vec{k})=\pm k\vec{\epsilon}^{\,(\pm)}(\vec{k})$ and $\vec{\epsilon}^{\,(\sigma)*}(\vec{k})\cdot\vec{\epsilon}^{\,(\sigma^{\prime})}(\vec{k})=\delta^{\sigma\sigma^{\prime}}$ with $k=|\vec{k}|$. Regarding quantization, we require the creation and annihilation operators, $a_{\vec{k}}^{(\sigma)}$ and $a_{\vec{k}}^{(\sigma)\dagger}$, to satisfy the usual canonical commutation relations, $[a^{(\sigma)}_{\vec{k}},a_{\vec{q}}^{(\sigma^{\prime})\dagger}]=\delta^{\sigma\sigma^{\prime}}\delta(\vec{k}-\vec{q})$.
@@
[…]
@@
Next, we discuss the effect when there is, e.g, a Dirac fermion $\psi$ interacting with gauge field during the inflation through $-g_{B-L}q_{\psi}\bar{\psi}\gamma^{\mu}X_{\mu}\psi$ where $q_{\psi}$ is the $B-L$ charge of this fermion. The amplified gauge fields induce the current of fermions, which backreacts on the dynamics of gauge fields. The extension to include many fermion species can be done by summing up their contributions. With one Dirac fermion species, the equation governing the energy transfer of the gauge field is given by
@@
which is derived from Eq. (II.6). Here, $\rho_{X}$ is the expectation value of the energy density of the gauge field, $\rho_{X}=(1/2\mathbb{V})\int d^{3}x\langle\vec{E}_{X}^{2}+\vec{B}_{X}^{2}\rangle$, and $\vec{J}_{\psi}$ is the induced current of fermion $\psi$. While the gauge fields evolve as stochastic variables Fujita et al. (2022a), they can be approximated as constant and anti-parallel electric and magnetic fields at the horizon scale, as previously discussed. In the presence of a constant magnetic field background, the spectrum of fermions forms Landau levels, with the lowest level exhibiting chiral asymmetry and the higher level exhibiting chiral symmetry. When electric fields are applied in parallel to magnetic fields, states are excited along the Landau levels, inducing a fermion current Nielsen and Ninomiya (1983); Fukushima et al. (2008). The induced current arises from contributions of the lowest Landau level, which describes the chiral anomaly Nielsen and Ninomiya (1983), and from the higher Landau levels, which account for the Schwinger effect Heisenberg and Euler (1936); Schwinger (1951), representing tunneling processes between the discrete energy levels. Collectively, by taking $\vec{E}_{X}=(0,0,E_{X})$ and $\vec{B}_{X}=(0,0,-B_{X})$ this current is evaluated as Domcke and Mukaida (2018)
@@
where the first factor of 3 outside big parenthesis accounts for 3 generations. The first and the second contributions come from up and down quarks of each color, while the remaining terms are from electrons and neutrinos. Here, for simplicity, we assume the $B-L$ charge of the right-handed neutrinos is $1$.
@@
denoting the effective instability parameter. The effect of fermion production is now implemented as a suppressed production of the helicity of the gauge field. We then estimate the resultant gauge field strength as follows. By assuming constant electromagnetic configuration, $\xi_{\rm eff}$ is also taken to be constant. The dynamical equilibrium equation is identical to the case where there are no fermions with $\xi\rightarrow\xi_{\rm eff}$ substituted in analytical solution, Eqs. (II.9) and (II.10). Given a specific value of $\xi$, equivalently $\dot{\phi}$, as it appears in Eq. (II.18), and assuming that the solutions for $E_{X}$ and $B_{X}$ take the same form as the analytical solutions in Eqs. (II.9) and (II.10), with $\xi$ replaced by $\xi_{\rm eff}$, one can determine the value of $\xi_{\rm eff}$ that satisfies the dynamical equilibrium condition, thereby obtaining the solution for the electromagnetic field configuration that includes the backreaction of fermions. The magnetic helicity at the end of inflation is also evaluated by Eq. (II.13) with $\xi$ replaced by $\xi_{\mathrm{eff}}$. As has been discussed, the produced fermions carry $B-L$ asymmetries. From Eq. (II.3), we evaluate the $B-L$ number density in the standard model sector at the end of inflation as

### v2

@@ Abstract
If the BLB-L symmetry is gauged with the addition of right-handed neutrinos, the standard model BLB-L current is anomalous with respect to the BLB-L gauge field itself. Then, the anomaly relation implies that the magnetic helicity of the BLB-L gauge field is related to the standard model BLB-L charges, although the whole universe is BLB-L neutral with right-handed neutrinos. Based on this, we propose a new leptogenesis scenario with the gauged BLB-L symmetry as follows. First, the magnetic helicity of the BLB-L gauge field is generated, e.g., by the axion inflation, together with the standard model and right-handed neutrino BLB-L charges, with the net BLB-L charge kept zero. The BLB-L charges in the standard model and right-handed neutrino sectors are then subject to washout effects from the interactions between them. After the washout effects decouple, the BLB-L gauge symmetry is Higgsed and the magnetic helicity of the BLB-L gauge field decays and generates BLB-L charges in the both sector; thanks to the washout effects, we obtain a non-zero BLB-L asymmetry. We show that the baryon asymmetry of the universe can be generated in this scenario, discussing the decay of the magnetic helicity of the BLB-L gauge field and the interactions between the right-handed neutrinos and the standard model particles.
@@ I Introduction
BLB-L charges are not washed out by the standard model process as the BLB-L symmetry in the standard model is not anomalous with respect to $\text{SU}(3)_{C}\times\text{SU}(2)_{L}\times\text{U}(1)_{Y}$. This, however, seems to conclude that BLB-L cannot be generated with the magnetic helicity, at first glance. Ref. Domcke et al. (2021, 2023) has circumvent the apparent contradiction by introducing the right-handed neutrinos, which explains the neutrino mass Workman et al. (2022) through the seesaw mechanism Minkowski (1977); Yanagida (1979); Gell-Mann et al. (1979). The focus of the work is on the conservation of fermion chiralities in the standard model; some of the standard model Yukawa couplings, such as those for $e,u$, and $d$, are small enough at high temperatures and their chiralities are effectively conserved in the early universe. The fermion chiralities are anomalous with respect to the standard model gauge groups and can be generated associated with the magnetic helicity. BLB-L violating interactions of the right-handed neutrinos, which become effective only when they are sufficiently massive before decay, can generate a non-zero BLB-L in the thermal bath with non-zero fermion chirality, even when the Sakharov conditions Sakharov (1967) are not simultaneously satisfied. Note that, however, this scenario depends on the flavor and chirality structure in the UV physics; many extensions of the standard model predict different flavor and chirality structures Isidori et al. (2010).
@@
In this work, we present a new avenue for the generation of the BLB-L asymmetry from magnetic helicity. Previous works assume that the BLB-L symmetry is not broken by the anomaly of the standard model gauge groups. However, this is not precise; the BLB-L symmetry has the gravitational anomaly and the ’t Hooft anomaly Weinberg (2013). The former is considered in the gravi-leptogenesis scenario Alexander et al. (2006), but, as Ref. Alexander et al. (2006) itself admits, the amount of the gravitational wave would be too much if we assumed all the baryon asymmetry of the universe is generated with helical gravitational waves. See also Refs. Fischler and Paban (2007); Kamada et al. (2020).
@@
On the other hand, the latter possibility has not been explored much, to our knowledge***See Ref. Chao et al. (2024) for a recent study to generate baryon asymmetry through the BLB-L ’t Hooft anomaly in axion inflation.. That is, once the BLB-L is gauged with the addition of right-handed neutrinos, the standard model BLB-L symmetry is anomalous with respect to the gauge field itself. Based on this, we propose a new mechanism for the generation of the baryon asymmetry of the universe as follows. The magnetic helicity of the BLB-L gauge field is generated in the early universe, e.g., during the axion inflation, and the BLB-L charges in the standard model sector are generated with the magnetic helicity. The BLB-L charges in the right-handed neutrino sector are also generated with the magnetic helicity, and the total amount of the BLB-L charges in the universe is zero. By the interactions between the right-handed neutrinos and the standard model particles, the BLB-L charges in both sectors once tend to be equilibrated and relaxed to zero while keeping the magnetic helicity of the BLB-L gauge field conserved. Then, as the BLB-L gauge symmetry is broken at the scale $v_{B-L}$, the magnetic helicity of the BLB-L gauge field decays and the difference between the BLB-L charges in the standard model sector and right-handed neutrino sector arises once more. If the washout interaction for the BLB-L charge in the standard model sector is decoupled before the decay of the magnetic helicity, the difference remains as the baryon number of the universe. As our scenario directly generates BLB-L charges, we may generate enough baryon asymmetry even if the flavor symmetry is badly broken in the UV physics.
@@
The organization of the paper is following. In Sec. II, we review the co-generation of the magnetic helicity and chirality and discuss the decay of the former. In Sec. III, we assume the interaction between the right-handed neutrinos and the standard model particles is such that the standard model neutrinos are Majorana particles and discuss the evolution of the BLB-L charges in the standard model sector. We find that the result depends on the UV structure of the theory and the phase transition of the BLB-L gauge symmetry breaking. In Sec. IV, we assume the standard model neutrinos are Dirac particles and discuss the evolution of the BLB-L charges in the standard model sector. We find that the baryon asymmetry of the universe can be easily generated in this case. Finally, Sec. V, is devoted to the discussion and conclusion.
@@ II Generation of the helicity and its decay
where $\vec{A}$ is the spatial component of the gauge field and $\vec{B}$ is the magnetic field. The magnetic helicity is a gauge invariant quantity if the magnetic field dumps off at the spatial infinity. In the following, we assume the BLB-L symmetry is gauged and consider the magnetic helicity of the BLB-L gauge field, $X_{\mu}$, namely, $\mathcal{H}_{X}$.
The magnetic helicity has a close relation to the chiral anomaly of the gauge field; the change of the magnetic helicity corresponds to the change of the chiral charge of the fermions. If we ignore the neutrino mass, the standard model BLB-L current, $J^{\mu}_{B-L,\text{SM}}$, is anomalous with respect to the BLB-L gauge fieldHere, we take the covariant form of the anomaly Bardeen and Zumino (1984).;
@@
where $g_{B-L}$ is the coupling constant of the BLB-L gauge field, $X_{\mu\nu}$ is the field strength tensor of $X_{\mu}$, and $\tilde{X}^{\mu\nu}$ is the dual tensor of $X_{\mu\nu}$. Integrating the above equation over the spatial volume, we obtain
@@
where $N_{B-L,\text{SM}}$ is the BLB-L charge of the standard model particles in the volume. Here, we assume that all fields drop off quickly enough at the spatial infinity. This equation shows that the change of the magnetic helicity of the BLB-L gauge field corresponds to the change of the BLB-L charge of the standard model particles.
In the following, we focus mainly on the dynamics of the BLB-L gauge field in the context of axion inflation. In the subsequent sections, we will analyze the evolution of the BLB-L asymmetry in the matter sector, taking into account the generation of BLB-L helicity in axion inflation. For this purpose, we need to specify the model of the right-handed neutrinos.
@@ II.1 Generation of the helicity
The Lagrangian of the inflaton $\phi$ and the BLB-L gauge field is given by
@@
where polarization vector $\vec{\epsilon}^{\,(\pm)}$ satisfying $\vec{\epsilon}^{\,(\pm)}(\vec{k})\cdot\vec{k}=0,\ i\vec{k}\times\vec{\epsilon}^{\,(\pm)}(\vec{k})=\pm k\vec{\epsilon}^{\,(\pm)}(\vec{k})$ and $\vec{\epsilon}^{\,(\sigma)*}(\vec{k})\cdot\vec{\epsilon}^{\,(\sigma^{\prime})}(\vec{k})=\delta^{\sigma\sigma^{\prime}}$ with $k=|\vec{k}|$. Regarding quantization, we require the creation and annihilation operators, $a_{\vec{k}}^{(\sigma)}$ and $a_{\vec{k}}^{(\sigma)\dagger}$, to satisfy the usual canonical commutation relations, $a^{(\sigma)}_{\vec{k}},a_{\vec{q}}^{(\sigma^{\prime})\dagger}]=\delta^{\sigma\sigma^{\prime}}\delta(\vec{k}-\vec{q})$.
@@
for sufficiently large $\xi$ ($\gtrsim 4$ Jiménez et al. (2017)), where $\langle\cdot\rangle$ denotes the quantum vacuum expectation value and the cutoff of the momentum is taken to be $2\xi aH_{\mathrm{inf}}$. See also Ref. Ballardini et al. (2019) for the renormalization of the gauge fields in axion inflation. We emphasize that the dominant contribution comes from superhorizon mode $k\lesssim aH_{\rm inf}$, so that the produced gauge field can be regarded to be constant over the Hubble scale, while the electric field and magnetic field point in an anti-parallel direction. The production rate of the physical helicity is given by
@@
Next, we discuss the effect when there is, e.g, a Dirac fermion $\psi$ interacting with gauge field during the inflation through $g_{B-L}q_{\psi}\bar{\psi}\gamma^{\mu}X_{\mu}\psi$ where $q_{\psi}$ is the BLB-L charge of this fermion. The amplified gauge fields induce the current of fermions, which backreacts on the dynamics of gauge fields. The extension to include many fermion species can be done by summing up their contributions. With one Dirac fermion species, the equation governing the energy transfer of the gauge field is given by
@@
which is derived from the equation of motion with the fermion current. Here, $\rho_{X}$ is the expectation value of the energy density of the gauge field, $\rho_{X}=(1/2\mathbb{V})\int d^{3}x\langle\vec{E}_{X}^{2}+\vec{B}_{X}^{2}\rangle$, and $\vec{J}_{\psi}$ is the induced current of fermion $\psi$. While the gauge fields evolve as stochastic variables Fujita et al. (2022a), they can be approximated as constant and anti-parallel electric and magnetic fields at the horizon scale, as previously discussed. In the presence of a constant magnetic field background, the spectrum of fermions forms Landau levels, with the lowest level exhibiting chiral asymmetry and the higher level exhibiting chiral symmetry. When electric fields are applied in parallel to magnetic fields, states are excited along the Landau levels, inducing a fermion current Nielsen and Ninomiya (1983); Fukushima et al. (2008). The induced current arises from contributions of the lowest Landau level, which describes the chiral anomaly Nielsen and Ninomiya (1983), and from the higher Landau levels, which account for the Schwinger effect Heisenberg and Euler (1936); Schwinger (1951), representing tunneling processes between the discrete energy levels. Collectively, by taking $\vec{E}_{X}=(0,0,E_{X})$ and $\vec{B}_{X}=(0,0,-B_{X})$ this current is evaluated as Domcke and Mukaida (2018)
@@
where the first factor of 3 outside big parenthesis accounts for 3 generations. The first and the second contributions come from up and down quarks of each color, while the remaining terms are from electrons and neutrinos. Here, for simplicity, we assume the BLB-L charge of the right-handed neutrinos is $1$.
@@
denoting the effective instability parameter. The effect of fermion production is now implemented as a suppressed production of the helicity of the gauge field. We then estimate the resultant gauge field strength as follows. By assuming constant electromagnetic configuration, $\xi_{\rm eff}$ is also taken to be constant. Let us assume that $\xi_{\rm eff}$ would be an independent and constant parameter. The equation of motion of $\vec{X}$ with the fermion current is identical to Eq. (II.6), where we do not have any fermion, with the replacement of $\xi\rightarrow\xi_{\rm eff}$. Therefore, the solutions for $E_{X}$ and $B_{X}$ would take the same form as the analytical solutions in Eqs. (II.9) and (II.10), with $\xi$ replaced by $\xi_{\rm eff}$. In reality, $\xi_{\rm eff}$ is a dependent and dynamical parameter, but once the system is in equilibrium and $\xi_{\rm eff}$ becomes constant, the values of $E_{X}$ and $B_{X}$ are expected to approach those values. With that solution for $E_{X}$ and $B_{X}$, the second term of the right-hand side of Eq. (II.19) is now a function of $\xi_{\rm eff}$ and Eq. (II.19) can be now regarded as a self-consistency equation on $\xi_{\rm eff}$ in terms of $\xi$. Thereby, we may estimate $\xi_{\rm eff}$, $E_{X}$ and $B_{X}$. This is called the “equilibrium estimate”. One may worry about the strong backreaction on the inflaton dynamics Garcia-Bellido et al. (2024); Figueroa et al. (2023); von Eckardstein et al. (2023); Figueroa et al. (2024), but if $\xi_{\mathrm{eff}}$ is small, regardless of the value of $\xi$, the backreaction is negligible and our treatment is consistent. The magnetic helicity at the end of inflation is also evaluated by Eq. (II.13) with $\xi$ replaced by $\xi_{\mathrm{eff}}$. Note that, there could be other constant solutions if $E_{X}$ and $B_{X}$ do not take the same form as Eqs. (II.9) and (II.10). In that case, the magnetic helicity becomes larger than the equilibrium estimate Gorbar et al. (2022) and the equilibrium estimate, which we adopt in this paper, is expected to be more conservative. All the estimates for the Schwinger backreaction thus far rely on the approximation that decomposes the electric and magnetic field in the expression of Schwinger current into the background and dynamical field, which is not very correct. To obtain a more precise estimate, numerical simulations with more appropriate treatment of the Schwinger current are needed. As has been discussed, the produced fermions carry BLB-L asymmetries. From Eq. (II.3), we evaluate the BLB-L number density in the standard model sector at the end of inflation as